OpenStudy (anonymous):

I'm working on derivatives and have encountered a notation I'm not familiar with. What derivative are you supposed to find if the dx/dy in front of the equation looks like d^2y/dx^2?

OpenStudy (anonymous):

\[\frac{ d^2y }{ dx^2 }\] Like that.

OpenStudy (calculusfunctions):

You have to find the second derivative (the derivative of the derivative).

OpenStudy (anonymous):

for all higher nth derivatives\[\frac{ d ^{n}y }{ dx ^{n} }=\frac{ d }{ dx }(y ^{(n-1)})\]The notation just says that this is the nth derivative, or the derivative of the n-1th derivative

OpenStudy (swag):

.